OpenStudy (anonymous):

find second derivative if y =(1-x)/(2-x)

OpenStudy (anonymous):

u know Quotient Rule..?

OpenStudy (anonymous):

i got 1st derivative as -1/(x^2-4x+4)...if thats even right...

zepdrix (zepdrix):

Yah I think that's correct. You don't need to expand out the bottom though, just leave it as a square.

OpenStudy (shamim):

\[\frac{ dy }{ dx }=\frac{ -1 }{ x ^{2}-4x+4 }\]

OpenStudy (shamim):

or,\[\frac{ d ^{2}y }{ dx ^{2} }=\frac{ (-1)(-1) (2x-4)}{(x ^{2}-4x+4)^{2} }\]

OpenStudy (shamim):

did u get my solution

OpenStudy (anonymous):

answer is (-2)/(2-x)^3 .......wish i know how to get that..

zepdrix (zepdrix):

\[\large y=\frac{(1-x)}{(2-x)}\]\[\large y'=\frac{(1-x)'(2-x)-(1-x)(2-x)'}{(2-x)^2}\] \[\large y'=\frac{(-1)(2-x)-(1-x)(-1)}{(2-x)^2} \quad =\quad \frac{-1}{(2-x)^2}\] \[\large y'=-(2-x)^{-2}\]Don't apply quotient rule for this next one, simply do power rule.\[\large y''=-(-2)(2-x)^{-3}(-1) \quad = \quad \frac{-2}{(2-x)^3}\]